OpenStudy (anonymous):

Please help? how do i solve (23pi/4)-pi ?

OpenStudy (anonymous):

\[\frac{ 23\pi }{ 4 } - \pi\]

OpenStudy (phi):

fyi, you can simplify or rewrite it. but "solve" means something else. think of this as adding two fractions or a fraction and a whole number. you need a "common denominator" once you have the same "bottom" for both, you add or subtract the tops, and keep the same bottom I would multiply the second number by 4/4 as the first step

OpenStudy (phi):

\[ \frac{ 23\pi }{ 4 } - \pi \\\frac{ 23\pi }{ 4 } - \frac{4\pi}{4}\] now subtract the tops \[ \frac{23\pi-4\pi}{4}\] up top you have 23 pi's take away 4 pi's , leaving 19 pi so the simplified form is \[ \frac{19\pi}{4}\]